Using Lagrange’s formula, find the values of

(i) y5 if y1 = 4, y3 = 120, y4 = 340, y5 = 2544

(ii) y0 if y–30 = 30, y–12 = 34, y3 = 38, y18 = 42.


1
Expert's answer
2021-01-03T16:01:24-0500

(i) y5 if y1 = 4, y3 = 120, y4 = 340, y5 = 2544

Answer y5=2544



y5 if y1 = 4, y3 = 120, y4 = 340, y6 = 2544

Here the intervals are unequal.


"\begin{matrix}
 x_0=1, & x_1=3, & x_2=4, & x_3=6 \\
 y_0=4, & y_1=120, & y_2=340, & y_3=2544
\end{matrix}"


By Lagrange’s interpolation formula we have

"y=f(x)=\dfrac{(x-x_1)(x-x_2)(x-x_3)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)}\times y_0"

"+\dfrac{(x-x_0)(x-x_2)(x-x_3)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)}\times y_1"

"+\dfrac{(x-x_0)(x-x_1)(x-x_3)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)}\times y_2"

"+\dfrac{(x-x_0)(x-x_1)(x-x_2)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)}\times y_3"



Put "x=5"

"f(5)=\dfrac{(5-3)(5-4)(5-6)}{(1-3)(1-4)(1-6)}\times 4""+\dfrac{(5-1)(5-4)(5-6)}{(3-1)(3-4)(3-6)}\times 120"

"+\dfrac{(5-1)(5-3)(5-6)}{(4-1)(4-3)(4-6)}\times 340"

"+\dfrac{(5-1)(5-3)(5-4)}{(6-1)(6-3)(6-4)}\times 2544"

"=1052"

"y5=1052"


(ii)

Here the intervals are unequal.


"\begin{matrix}
 x_0=-30, & x_1=-12, & x_2=3, & x_3=18 \\
 y_0=30, & y_1=34, & y_2=38, & y_3=42
\end{matrix}"


By Lagrange’s interpolation formula we have

"y=f(x)=\dfrac{(x-x_1)(x-x_2)(x-x_3)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)}\times y_0"

"+\dfrac{(x-x_0)(x-x_2)(x-x_3)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)}\times y_1"

"+\dfrac{(x-x_0)(x-x_1)(x-x_3)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)}\times y_2"

"+\dfrac{(x-x_0)(x-x_1)(x-x_2)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)}\times y_3"



Put "x=0"

"f(0)=\dfrac{(0+12)(0-3)(0-18)}{(-30+12)(-30-3)(-30-18)}\times 30"

"+\dfrac{(0+30)(0-3)(0-18)}{(-12+30)(-12-3)(-12-18)}\times 34"

"+\dfrac{(0+30)(0+12)(0-18)}{(3+30)(3+12)(3-18)}\times 38"

"+\dfrac{(0+30)(0+12)(0-3)}{(18+30)(18+12)(18-3)}\times 42"

"=\dfrac{409}{11}"

"y_0=\dfrac{409}{11}"